### v1

# Think Before You Diffuse: LLMs-Guided Physics-Aware Video Generation

Abstract
Recent video diffusion models have demonstrated their great capability in generating visually-pleasing results, while synthesizing the correct physical effects in generated videos remains challenging. The complexity of real-world motions, interactions, and dynamics introduce great difficulties when learning physics from data. In this work, we propose DiffPhy, a generic framework that enables physically-correct and photo-realistic video generation by fine-tuning a pre-trained video diffusion model. Our method leverages large language models (LLMs) to explicitly reason a comprehensive physical context from the text prompt and use it to guide the generation. To incorporate physical context into the diffusion model, we leverage a Multimodal large language model (MLLM) as a supervisory signal and introduce a set of novel training objectives that jointly enforce physical correctness and semantic consistency with the input text. We also establish a high-quality physical video dataset containing diverse phyiscal actions and events to facilitate effective finetuning. Extensive experiments on public benchmarks demonstrate that DiffPhy is able to produce state-of-the-art results across diverse physics-related scenarios. Our project page is available at https://bwgzk-keke.github.io/DiffPhy/
Figure 1: DiffPhy enables physically grounded and semantically aligned video generation across diverse real-world scenarios, including gravity-driven motion, fluid interactions, forceful impacts, and object manipulation. It outperforms state-of-the-art video diffusion model Wan 2.1–14B [50] in both visual plausibility and physical coherence.
1 Introduction
Recent advances in video diffusion models have made it possible to generate high-fidelity videos with rich details from text prompts [15, 50, 63, 10, 27, 2, 40, 35, 52, 13, 42, 45, 26]. Such generation capability has opened up a wide range of real-world applications, including movie production [65], gaming [48], and even serving as virtual environments for robotics and embodied AI learning [2, 3]. Despite the great visual quality, accurately simulating the correct physical-related effects remains challenging and is often overlooked by existing methods [25, 61]. These models learn physics by exhaustively training the model over a large corpus of video data. The complexity of real-world motion, interactions, and dynamics make such implicit learning difficult and less effective [7].
In contrast, traditional graphics approaches [19, 20, 21, 39, 34] predefine physical parameters for a scene using simulation engines, enabling them to render results that are inherently physically accurate. However, manually specifying these physical properties is often labor-intensive and becomes impractical as scene complexity increases. As a result, such methods are typically constrained to simple physical scenarios, such as perfectly elastic collisions [25] and basic rigid body dynamics [34]. The capbility to scale to complex real-world environments, where physical effects emerge from intricate motion, interactions, forces, collisions, and environmental context, remains an open challenge.
In this work, we propose DiffPhy, a novel framework that enables video diffusion models to generate physically accurate and visually compelling videos from arbitrary user prompts. At the heart of our approach is the use of Large Language Models (LLMs) [1, 7] as both reasoners and evaluators to guide the generation process. Specifically, our method begins by using an LLM [1] to infer a rich physical context from the input text. This inferred context is then integrated into the prompt, yielding an enhanced text instruction that provides explicit physical cues for video generation. We then fine-tune a pre-trained video diffusion model [50] using this enhanced prompt to make it physics-aware. During fine-tuning, a Multimodal Large Language Model (MLLM) [7] evaluates the generated video by analyzing their physical alignment and semantic consistency with the text description as well as their physical plausibility. This supervision is implemented as a set of novel training objectives that enforce both physical correctness and semantic consistency with the original prompt.
One challenge left is how to obtain a high-quality, annotated video dataset that captures diverse physical scenarios for effective fine-tuning. Most existing datasets are designed primarily for evaluation [6, 7, 34, 38], and typically consist of synthetic or simulated scenes with limited scale, diversity, and complexity. Fine-tuning directly on such datasets often leads to degraded performance. To this end, we also introduce a larger curated dataset of real videos that covers diverse entities, motions, behaviors and physical phenomena. This dataset can be effectively used for training video generation model, boosting their capabilities to handle complex real-world physical scenarios.
Our contribution can be summarized in fourfold: (i) We introduce DiffPhy, the first framework that enables text-to-video diffusion models to generate physically accurate and visually realistic videos from arbitrary prompts. (ii) We propose a novel training strategy that uses LLMs for physical reasoning and MLLMs for supervision, enforcing both physical correctness and semantic alignment during fine-tuning. (iii) We introduce a curated dataset of real-world videos covering diverse physical scenarios to support effective training for physics-aware video generation. (iv) Our DiffPhy achieves state-of-the-art performance in generating physically plausible and semantically coherent videos across diverse scenarios.
2 Related Work
Text-to-Video Generation. Diffusion-based text-to-video generation is undergoing a rapid revolution, with improvements in quality [2, 63, 8], controllability [67, 4, 44], efficiency [58, 66, 24], extrapolation [68, 11, 36], etc. Incorporating pre-trained text-to-image diffusion models, early works, such as AnimateDiff [15] and SVD [8], modify the UNet blocks to model the temporal dependencies. Recent advancements [63, 27, 51] leverage the scalable Diffusion Transformer [41] (DiT) with more parameters and training data and present better synthesis quality and diversity. Closed-source models, such as SORA [40], Pika [42], and Kling [26], have demonstrated their impressive capabilities in video production. However, these existing methods implicitly learn the physical principles from data distribution and often neglect real-world physical laws [7], focusing more on visual aspects like color and texture rather than underlying physical dynamics [25]. Consequently, the generated videos frequently suffer from a lack of physical commonsense and fail to capture authentic object motion.
Video Physics Reasoning. Existing work has used physics simulators and engines for tasks such as rigid body dynamics, fluid simulation, and deformable object modeling [19, 20, 21, 39, 34]. Recent methods [12, 29, 30, 56, 55, 59, 60] extract physical representations from visual inputs via neural networks for downstream reasoning or simulation. While effective in domains like graphics, robotics, and scientific computing, these approaches often rely on rule-based, handcrafted solvers, which limits their expressiveness, scalability, and requiring extensive manual tuning. Moreover, they are typically confined to synthetic data and can not scale to complex, real-world video generation.
LLM in T2V Generation. LLMs have been widely used for prompt refinement in text-to-image and text-to-video generation, helping interpret prompts and guide layout planning [9, 22, 31, 33, 57, 62, 69]. Methods like PhyT2V [61] refine prompts iteratively, but they operate solely in the text domain and do not improve the generation model itself. Recent work has proposed physical evaluation benchmarks that assess physical plausibility using multimodal LLMs [7, 6, 37], leveraging models such as VideoCon [5], GPT-4o [1], LLaVA [32], and InternVideo [53]. We adopt the multimodal LLM evaluator VideoCon-Physics [7] for its lightweight, open-source design. It is built on VideoCon [5], with CLIP [43] as the visual encoder and LLaMA-7B [47] as the language backbone.
3 Method
Figure 2: An overview of DiffPhy. Our method contains five steps. 1. Given a user prompt, we leverage a pretrained LLM to reason physical properties from the text input. 2. We then (a) enhance the user prompt with physical context and (b) produce a list of relevant physical phenomena associated with the described event. 3. We use the enhanced prompt to guide video generation. The phenomena list is used to penalize outputs with implausible physics. 4. A set of novel training objectives are proposed to jointly enforce physical correctness and semantic consistency.
We present DiffPhy, a generic framework that enables diffusion models to generate physically-correct and photo-realistic videos. Our core idea is to (1) ground the user prompt with rich physical context through LLM reasoning and (2) fine-tuning the diffusion model to make it physics-aware by ensuring the generated physical effects are valid. An overview of our method is shown in Figure 2. In more detail, given a user prompt, we first apply a pretrained LLM (e.g. GPT-4) to reason the underlying physical attributes (e.g. gravity, velocity, collisions, bounce) via Chain-of-Thought [54] prompting (Figure 2-(1)). Then, based on the inferred properties, we query the LLM again to enhance text prompt to include physical context (Figure 2-(2a)). The enhanced text description is then fed to the video diffusion model to guide the generation process (Figure 2-(3)). We also request the LLM to generate a list of physical phenomena (Figure 2-(2b)) associated with the described event. During training, this list is used to penalize any synthesized videos that contain invalid physical effects (Figure 2-(4)). We make the diffusion model physics-aware by training with three novel loss terms that jointly encourage the physical correctness and video-realism. In the following, we first describe how to obtain enhanced text prompt with detailed physical attributes (Section 3.1) and then discuss how to train the diffusion model to produce physically-correct videos (Section 3.2).
3.1 Grounding Text Description with Physical Context
As shown in Figure 2, user prompts are often simple and incomplete, without detailed descriptions about the physical effects associated with the event/action. Existing methods [50, 40, 49, 63, 8] implicitly interpret physics from such text by exhaustively training over a large data corpus, which is known to be difficult and costly [7]. Instead, our method seeks to explicitly mine the physical attributes/effects in the user prompt as a pre-processing step. Thereafter, the generation process can be guided by an enhanced text prompt that contains sufficient physical context. Specifically, we leverage a pre-trained LLM to reason physical context and obtain three types of text descriptions: (1) a list of physical attributes that capture roles (cause/effect), interactions, and physics; (2) an enhanced prompt with scene and physical details; (3) a list of physical phenomena associated with the event.
Physical attributes reasoning.
To extract physical attributes from text, we prompt a pretrained LLM using chain-of-thought (CoT) [54]. Specifically, we prompt the language model by “list all relevant physical principles, explain their roles, and formalize them as quantitative constraints." This allow the LLM to effectively identify event semantics including underlying forces (e.g., gravity, friction), kinematic relationships (e.g., constant velocity, acceleration), and interaction rules (e.g., elastic vs. inelastic collisions). We then parse the LLM output into structured machine-readable rules, creating a coherent physical description of the target scene. In practice, the user prompt might not be complete and can sometime omit important entities or interactions. For example, an incomplete text prompt might like “A balloon is popped by being over-inflated, causing it to burst", which does not specify the cause for over-inflation. As a result, a balloon that pops without cause becomes an invalid scene. To address this, we extend the CoT reasoning into a procedural process through four detailed steps: (1) identifying the key physical principles involved; (2) determining which entities initiate actions and which are affected; (3) reasoning about how these entities interact; and (4) noting any resulting physical phenomena. As a result, we obtain a list of physical attributes that are both comprehensive and valid. The detailed prompt is provided in Appendix.
Enhanced text prompt.
Building on the extracted physical attributes, we prompt the LLM to “vividly describe the scene, including the objects, actions, and atmosphere." This generates a richer, physics-aware text description that captures not only the mechanical aspects but also the narrative depth of the scene. We allow LLM to introduce new entities for logical consistency. For example, if a user prompt describes “a candle falls" in Figure 2, the LLM will also hallucinate an external force (e.g., from a cat) to make the event physically plausible (Figure 2-2(a)). The resulting text prompt, enhanced with physical context and narrative clarity, is then used as input to guide video generation.
Physical phenomena.
We also interested in obtaining a list of physical phenomena associated with the target event. These physical phenomena are facts caused by the changes of the event state, such as an action of an entity and a consequence of a physical law. In case of "a candle falling to its side", one fact might be "The candle’s flame is taller when burning intensely" and "changes position from upright to lying down". The full prompt used to obtain such list is included in Appendix. These physical phenomena depict the facts to be expected in the generated videos and therefore can be used to guide the training process. For example, one can reward diffusion model when the fact is met and penalize it when the fact is missed or wrong.
3.2 Physics-aware Model Training
One challenge left is how to train a diffusion model to effectively leverage the rich physical context in the text description as well as precisely generate the desired effects. This indicates that the diffusion model needs to (1) comply with specific physical rules implied in the text prompt; (2) respect the commonsense as a natural video and (3) produce desired content adhering to the text. Ideally, this can be accomplished by optimizing the diffusion model towards some metrics that evaluate how well a generated video is aligned with specific physical effects. Yet to the best of our knowledge, there is no such measurement in the literature. To bridge this gap, we propose to adopt a multi-modal LLM (MLLM) [7] as a learned evaluator, learning its strong capability to perform complex physical reasoning over video and text. Specifically, We design three MLLM-based losses, i.e., physical phenomena loss, commonsense loss, semantic consistency loss, alongside the standard denoising score matching objective [16]. These losses are computed by querying MLLM over generated videos.
Physical phenomena loss.
It is essential to make sure that the diffusion model follows the physical principles when generating a target action/event. This is conceptually difficult to verify as the physical rules are abstract. To overcome this, we instead validate whether the generated results contain the desired facts listed in the physical phenomena, as a way to indirectly request the diffusion model to respect the underlying physical principles.
Specifically, during training, at a sampled timestep $t$, we decode the predicted latent clip $x_{t}\in R^{m\times c\times h\times w}$ with $m$ frames into a pixel space video $v_{t}$. Then, we employ an MLLM [7] to evaluate the alignment between $v_{t}$ and each fact $f_{i}$ in the physical phenomena $\mathbf{F}=\{f_{1},...,f_{n}\}$, one by one. For each fact $f_{i}$, the MLLM outputs an integer label $l_{i}\in\{0,1\}$, indicating the fact is ‘"not matched" or "matched" in $v_{t}$. This allows us to define a loss function to encourage the diffusion model to produce "matched" results and penalize "unmatched" results. Specifically, for a fact $f_{i}$, we can compute an expected label $\mathbb{E}(l_{i})$ by weighted averaging over all labels. Formally,
$\mathbb{E}(l_{i})=\sum_{l\in\{0,1\}}l\cdot p(l),\quad\text{where the probability }p(l)=\frac{e^{z_{l}}}{\sum_{k\in\{1,2\}}e^{z_{k}}}$ (1)
where $z_{k}$ corresponds the logit associated with the word $k$ predicted by the MLLM. Given the expected label, we then define the loss function as $\mathcal{L_{\text{phen}}}=\sum_{f_{i}\in\mathbf{F}}||\mathbb{E}(l_{i})-1||_{2}^{2}$. Training the model with $\mathcal{L_{\text{phen}}}$ encourages the diffusion model to generate the desired physical effects.
Resolving failure facts via attention injection.
It is equally important for the diffusion model to resolve any failed or incorrect facts. For a "not matched" fact $f_{i}$, we introduce an additional module to guide the diffusion model to better understand the failure or missed cases during generation. This module encodes $f_{i}$ into text embeddings and conditions the diffusion model via a separate cross-attention layer in each DIT block, as shown in Figure 2-(3). To stablize the training, we copy the original cross-attention module and connect it with the original model via a zero-initialized convolution layer (ZeroConv) in a way similar to [64]. To reduce training cost, we apply a LoRA adapter [18] to the cloned attention module and fine-tune only the adapter. More implementation details can be found in the Appendix.
Physical commonsense loss.
We also introduce a physical commonsense loss $\mathcal{L}_{\text{com}}$ to evaluate the overall physical plausibility of the scene. We use the same multi-modal LLM [7] again to assign an integer-valued plausibility score $s_{com}$ to $v_{t}$, ranging from 1 to 5, where 5 indicates the highest plausibility. Similar to $\mathcal{L}_{phen}$, we compute an expected score $s_{\text{com}}$ over all possible values (i.e. {1,2,3,4,5}) using Eq. 1. Then, the physical commonsense loss can be defined as a mean squared error (MSE): $\mathcal{L}_{com}=\|\mathbb{E}(s_{\text{com}})/\tau-1\|_{2}^{2}$, where $\tau=5$ is a normalization constant. This loss encourages the diffusion model to generate videos that meet physical commonsense.
Semantic consistency loss.
To ensure the generated content still follows the input prompt, we introduce a semantic consistency loss. Again, we use the same pretrained multi-modal LLM [7] to give scores from 1 to 5 for evaluating the semantic consistency between $v_{t}$ and the input text prompt. The semantic consistency loss has the same format as physical commonsense loss, i.e. $\mathcal{L}_{sem}=\|\mathbb{E}(s_{\text{sem}})/\tau-1\|_{2}^{2}$, where $\tau=5$ is a normalization constant, $s_{\text{sem}}$ is the expected score over all possible values (i.e. {1,2,3,4,5}).
Training objective.
The final loss is computed by combining the original denoising score matching $\mathcal{L}_{\epsilon}$ objective [16] with the three newly proposed MLLM-based loss functions, i.e., phenomena loss $\mathcal{L}_{\text{phen}}$, physical commensense loss $\mathcal{L}_{\text{com}}$ and semantic consistency loss $\mathcal{L}_{\text{sem}}$. Formally,
$\mathcal{L}=\mathcal{L}_{\epsilon}+\beta(\mathcal{L}_{\text{phen}}+\mathcal{L}_{\text{com}}+\mathcal{L}_{\text{sem}})$ (2)
where $\beta$ is a balancing hyperparameter.
Training details. As shown in Figure 2, we freeze the original DIT blocks and train only LoRA layers to reduce training cost. When a failure fact occurs, we sample the same diffusion step again with the failure fact-based guidance injected. We also randomly perform a training step without activating the injection branch. This allows diffusion model to still generate high-quality videos from solely text input. More implementation details are provided in Appendix.
Figure 3: Qualitative comparison with T2V models on the VideoPhy2. We show two challenging cases, i.e., sports and box collapse, where our results are more natural and description-consistent.
4 Experiments
HQ-Phy Dataset. We curate a dataset of approximately 8,000 videos selected from the VIDGEN-1M dataset [46], which includes the video, caption, enhanced prompts, and extracted physical phenomena labels. We filter the dataset to exclude cartoon and other unrealistic scenes. To extract the enhanced prompts and detailed physical phenomena from this training dataset, we use DeepSeek-R1-32B [14] as a cost-effective alternative to GPT. Details are provided in Appendix.
Evaluation benchmarks: In our experiments, we utilize two datasets specifically designed to evaluate physical commonsense reasoning in text-to-video generation. VideoPhy2 [7] provides a collection of 590 human-verified captions that describe realistic interactions between various physical entities, including both solid and fluid objects. This dataset serves as a testbed for assessing whether generated videos reflect plausible real-world physics. Additionally, we incorporate PhyGenBench [37], a benchmark constructed to further probe the physical awareness of T2V models. It comprises 160 designed prompts spanning four fundamental domains of physics: mechanics, optics, thermal processes, and material properties.
{NiceTabular}
l|cccccc Model Physical domains($\uparrow$) Average
Mechanics Optics Thermal Material
CogVideoX-2B 0.38 0.43 0.34 0.39 0.37
CogVideoX-5B 0.43 0.55 0.40 0.42 0.45
Open-Sora V1.2 0.43 0.50 0.44 0.37 0.44
Lavie 0.40 0.44 0.38 0.32 0.36
Vchitect 2.0 0.41 0.56 0.44 0.37 0.45
Pika 0.35 0.56 0.43 0.39 0.44
Kling 0.45 0.58 0.50 0.40 0.49
Wan 0.36 0.53 0.36 0.33 0.40
DiffPhy (ours) 0.53 0.59 0.58 0.46 0.54
Table 1: T2V comparisons on PhyGenBench.
Figure 4: Visualization of Table 1.
{NiceTabular}
l|cccccc Dimension Methods Physical metrics ($\uparrow$) Overall Quality Average
Phenomena Order GPT-4o Open
Mechanics Kling 0.608 0.454 0.533 0.508 0.450
Wan 0.608 0.400 0.442 0.483 0.358
DiffPhy 0.733 0.525 0.617 0.567 0.533
Optics Kling 0.753 0.550 0.660 0.660 0.580
Wan 0.760 0.563 0.587 0.573 0.527
DiffPhy 0.833 0.660 0.667 0.647 0.587
Thermal Kling 0.611 0.372 0.544 0.567 0.500
Wan 0.500 0.272 0.378 0.500 0.356
DiffPhy 0.700 0.578 0.622 0.656 0.578
Material Kling 0.617 0.342 0.492 0.483 0.400
Wan 0.575 0.254 0.392 0.417 0.333
DiffPhy 0.725 0.433 0.575 0.500 0.458
Table 2: PhyGenBench evaluation of phenomena detection, physical order, GPT-4o and open-source models.
Figure 5: Average metrics of Table 2.
Metrics. To ensure a comprehensive evaluation of our generated videos, we adopt a multi-faceted assessment strategy combining model-based metrics and human judgment. Firstly, we adopt the evaluation metrics from the PhyGenBench benchmark [37]. Following PhyGenBench [37], we report key phenomena detection, physical order verification, and average scores to indicate overall quality using GPT-4o [1] and open-source models, i.e., CLIP [43], InternVideo2 [53], and LLaVA [32]. Second, we use the VideoCon-Physics evaluator from VideoPhy2[7], which reports two metrics: physical commonsense (PC) and semantic adherence (SA), each scored from 1 to 5. PC evaluates physical plausibility, while SA measures alignment with the input prompt. Following VideoPhy2[7], we set the overall score to 1 if both PC and SA are $\geq$ 4, and 0 otherwise. Finally, we conduct a gold-standard human evaluation via blind review, where three annotators independently score each video on semantic adherence and physical commonsense. This complements automatic metrics with qualitative insight.
Baselines. We apply DiffPhy on Wan2.1-14B [50], a DiT-based open-source T2V models, and compare it to eleven open or close-source T2V models, including Wan2.1-14B [50], CogVideoX-5B [63], VideoCrafter2 [10], HunyuanVideo-13B [27], Cosmos-Diffusion-7B [2], OpenAI Sora  [40], Luma Ray2 [35], LaVie [52], Vchitect [13], Pika [42], and Kling [26]. Although DiffPhy supports various backbones, we use Wan2.1-14B for illustration.
Figure 6: Visualization of (a) Open-source T2V model comparison on physical activities, and (b) ablation study on physical interactions. All prompts and compared videos are from VideoPhy2.
4.1 PhyGenBench Evaluation
Table 1 and Figure 4 present the performance of our model compared with a range of existing methods, including CogVideoX (2B/5B) [63], OpenSora [40], LaVie [52], Vichitect [13], Pika [42], Kling [26], and Wan 2.1-14B [50]. The results reported in Table 1 are sourced from the official PhyGenBench leaderboard [37], with the results for Wan 2.1-14B reproduced using its publicly released code and model [50]. Our DiffPhy not only outperforms Wan by a substantial margin across all metrics but also achieves the highest average score and the best performance in all of the dimensions, demonstrating the effectiveness of our proposed strategies.
Physical phenomena: We further validate the ability of our method to capture detailed physical phenomena. Table 2 and Figure 5 compare the performance of our model, the baseline Wan, and the second-ranked closed-source model Kling [26] on aspects including key phenomena (phenomena), physical order (order), and overall quality. The overall quality is indicated by the average results from GPT-4o [1] and open-sourced models (Open), i.e., CLIP [43], LLaVA [32], and InternVideo2 [53]. Our model achieves the highest scores in both key phenomena and physical order, with strong average performance across four physics-related categories: Mechanics, optics, thermal, and material properties. These results indicate that our model not only identifies key physical phenomena accurately but also generates videos that follow coherent physical progression.
Table 3: Comparison of model-based and human-based scores on the VideoPhy2 dataset.
{NiceTabular}
lcccccccc Methods Model-based Human-based
Hard Activity Interaction Overall Hard Activity Interaction Overall
Videocrafter 0.054 0.201 0.229 0.205 0.131 0.101 0.131 0.105
Sora 0.053 0.244 0.200 0.217 0.267 0.222 0.267 0.233
Ray2 0.067 0.259 0.204 0.234 0.185 0.210 0.185 0.203
Hunyuan 0.096 0.255 0.312 0.267 0.159 0.176 0.159 0.172
Cosmos 0.034 0.219 0.223 0.229 0.274 0.226 0.274 0.241
cogvideo 0.051 0.258 0.236 0.261 0.000 0.246 0.261 0.250
Wan 0.034 0.210 0.255 0.225 0.219 0.315 0.362 0.326
DiffPhy (w/o CoT) 0.045 0.245 0.255 0.247 0.386 0.519 0.400 0.489
DiffPhy (w/ CoT) 0.084 0.266 0.317 0.281 0.421 0.526 0.556 0.533
4.2 Videophy bechmark
Table 4.1 and Figure 3 demonstrate that our method DiffPhy, significantly outperforms the existing text-to-video (T2V) approaches on the VideoPhy2 dataset. Compared to Wan, our model shows notable improvements in average performance across all subcategories, including sports and physical activities (Activity), object interactions (Interaction), and challenging physical scenarios (Hard). We evaluate our method on the VideoPhy2 benchmark, following the protocol outlined in [7]. To ensure a fair comparison, we report results both with and without CoT prompting. While our baseline already surpasses Wan, incorporating CoT further boosts the performance, achieving the highest average scores across all video categories. Figure 6 visualizes typical videos generated by our model with and without CoT reasoning, compared to the baseline Wan model. This demonstrates DiffPhy ’s ability to generate coherent and plausible videos. We also observe discrepancies between model-based and human-based ratings, which is also mentioned in [7]. This discrepancy arises because although model-based metrics capture some aspects of physical correctness, human evaluation remains the definitive standard for assessing physical plausibility. Results for the compared methods, including both model and human evaluations, are taken from the original VideoPhy2 leaderboard [7]. Our own human evaluation results are provided in Table 4.2.
Human evaluation. We conduct a human evaluation on two dimensions consistent with VideoPhy2 [7], including Semantic Alignment (SA) and Physical Commonsense (PC). Table 4.2 compares our model with the top two models on the Videophy2 human-annotated leaderboard, CogVideoX [17] and Wan 2.1-14B [51]. Our model outperforms Wan across all three evaluation dimensions, even without Chain-of-Thought (CoT). CoT further boosts performance across all dimensions.
Table 4: Human evaluation on VideoPhy2 dataset.
{NiceTabular}
lcccccccccccc Methods Hard Activity Interaction Average
SAPCOverallSAPCOverallSAPCOverallSAPCOverall
CogVideoX2.018 2.193 0.018 2.452 2.415 0.141 2.311 2.467 0.022 2.417 2.428 0.111
Wan3.439 3.421 0.298 3.489 3.504 0.378 3.578 3.378 0.311 3.511 3.472 0.361
DiffPhy (w/o CoT) 3.456 3.439 0.386 3.793 3.704 0.519 3.556 3.489 0.400 3.733 3.650 0.489
DiffPhy (w/ CoT) 3.702 3.491 0.421 3.800 3.770 0.526 3.867 3.822 0.556 3.817 3.783 0.533
Table 5: Ablation results on VideoPhy2. $\mathcal{L}_{\text{com}}$, $\mathcal{L}_{\text{sem}}$,$\mathcal{L}_{\text{phen}}$ denote physical commonsense loss, semantic consistency loss, and physical phenomena loss, respectively; Iterative refers to an optional second-pass refinement by injecting the failure phenomena during inference.
{NiceTabular}
cccccccccccc Method Training Inference Model-based Human-based
$\mathcal{L}_{\text{com}}$&$\mathcal{L}_{\text{sem}}$ $\mathcal{L}_{\text{phen}}$ Enhance Prompt IterativeActivityInteractionOverallActivityInteractionOverall
Wan $\times$ $\times$ $\times$ $\times$ 0.210 0.255 0.225 0.361 0.378 0.311
#1 $\checkmark$ $\times$ $\times$ $\times$ 0.238 0.267 0.245 0.382 0.441 0.386
#2 $\checkmark$ $\checkmark$ $\times$ $\times$ 0.245 0.255 0.247 0.400 0.519 0.489
#3 $\checkmark$ $\times$ $\checkmark$ $\times$ 0.259 0.276 0.270 0.478 0.523 0.492
#4 $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 0.266 0.317 0.281 0.526 0.556 0.533
#5 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.273 0.323 0.289 0.535 0.560 0.542
4.3 Ablations
We conduct ablation studies to evaluate the effectiveness of each design choices. Results are reported in Table 4.2. We have the following findings. First, incorporating MLLM-based loss functions yields substantial gains over the Wan [50] baseline on both automatic and human evaluation. Second, adding the attention injection to resolve failure facts with phenomena loss ($\mathcal{L}_{\text{phen}}$) further improves performance. Third, incorporating chain-of-thought (CoT) prompting enhances prompt clarity and reasoning, resulting in additional improvement. Visual results on the effects of CoT are reported in Figure 6. During inference, our DiffPhy framework supports an optional two-pass inference by correcting failure cases from previous pass via attention injection, which yields additional performance improvements. Because it introduces runtime overhead, for fair comparison, we only report results using one-pass inference for the main experiments.
5 Conclusion
We introduce DiffPhy, a novel framework that enables diffusion models to generate physically-correct and realistic videos from arbitrary user prompts. This is achieved by leveraging Large Language Models (LLMs) to reason physical context from the input prompt and use it to guide the diffusion process. We then finetune the diffusion model to make use of this context using a Multimodal LLMs (MLLMs) as supervisory signals. Additionally, we provide a curated dataset of real-world videos covering a wide range of physical phenomena. Experiments on public benchmarks demonstrate that DiffPhy achieves state-of-the-art performance for physics-aware video generation.
Appendix A Video Demonstration
We encourage readers to watch the supplemental video, which showcases visual results and comparisons, providing a more comprehensive illustration of both the physical realism and generation quality of DiffPhy.
Appendix B Implementation Details
Experiment setup:
All experiments are conducted using four NVIDIA H100 GPUs with a batch size of 4. We fine-tune all models for 10 epochs, using Wan 2.1-14B [50] as the base model, following its training paradigm and data preprocessing setup. Specifically, we fine-tune the LoRA adapter on top of the Wan 2.1-14B backbone, using a LoRA alpha of 32 and a rank of 16. The hyperparameter $\beta$ of the training objective is empirically set to 0.1. To maintain training flexibility, we enable the diffusion model to occasionally skip applying the physical phenomena violation guidance, with a probability of 0.1. Thus, the model can generate high-quality videos even without the guidance of physical phenomena.
LLM:
We use LLM to perform Chain-of-Thought (CoT) reasoning, to generate enhanced prompt and physical phenomena. Specifically, we employ GPT O4-mini [23] for CoT during inference. For the large-scale training dataset, we use DeepSeek-R1-32B [14] as a cost-effective open-source option. The detailed prompts are provided in  C.
MLLM:
[…]
Evaluation prompts:
When generating videos for benchmark evaluation, i.e., PhyGenBench and VideoPhy, we use the same input prompts across all compared methods and DiffPhy (w/o CoT). These prompts are taken directly from the benchmarks provided by VideoPhy2 [7] and PhyGenBench [37], using their augmented prompts of the original user inputs. For DiffPhy (w/ CoT), we use our own enhanced prompts generated through Chain-of-thought physical reasoning.
Human evaluation:
We use human evaluation results as the ground-truth metric to complement model-based evaluations. We design a blind evaluation platform and conduct a blind review process for human annotation, where each annotator is asked to assess the physical commonsense and semantic alignment of text-to-video generation results. To ensure consistency, we follow the same video sampling strategy used by [7] for evaluating Sora.
Curation of PhyHQ dataset:
We curate the PhyHQ dataset to enable effective finetuning of physics-aware video diffusion models. Firstly, we collect the videos selected from the VIDGEN-1M dataset [46], which includes the videos and corresponding captions. Then, we further extract physical phenomena labels using the open-source DeepSeek-R1-32B [14], guided by the prompt described in Section C. This serves as a cost-effective alternative to using GPT-based models. We filter out the unrealistic videos with zero-shot-classification of BART-large-mnli [28] based on the video descriptions. For dataset preprocessing, we follow the sampling and process pipeline of Wan 2.1  [50]
Appendix C Chain-of-thought (CoT)
C.1 CoT Generation
During the Chain-of-Thought (CoT) reasoning process, the LLM is instructed to identify relevant physical laws, the resulting phenomena, the affected entities, and how their behavior changes as a consequence. To clarify the expected output structure, we leverage in-context learning—a technique where models are provided with annotated examples within the prompt to illustrate the desired reasoning and response format—ensuring the LLM aligns its output with task-specific requirements. To maintain coherence and minimize noise in the generated responses, we instruct the LLM to consistently reuse terminology from the original prompt when referring to the same entities. We also instruct the LLM to focus more on describing observable phenomena rather than abstract physical laws in the enhanced prompt, as phenomena are generally easier for the diffusion model to interpret and learn from than explicit rules. Additionally, to avoid overwhelming the model with excessive context, which can obscure key interactions, we explicitly limit the length of the enhanced prompt in the instructions. The prompt used for CoT generation is shown below.
Instruction for CoT Generation Analyze the scene by: 1. Identifying the key physical principles at work. 2. Determining which entities initiate actions and which are affected. 3. Examining how these entities interact. 4. Noting any observable physical phenomena that result. 5. Describe the scene vividly and clearly, capturing the key actions, objects, and atmosphere in natural language. Ensure that all objects mentioned in the original caption are included in the output, using the same wording for each entity. Also, include any entities or individuals involved in the interaction to ensure the scenario remains logical and coherent. The output should be a natural and simple description of the observed phenomena, without the detailed physical analysis, written in plain language and limited to 70 words. Input scene: A delivery drone descends into a residential courtyard, avoiding tree branches and landing precisely on a designated pad. Output enhanced prompt: A compact delivery drone hovers quietly above a residential courtyard, its rotors emitting a soft hum. It skillfully avoids low-hanging branches and lands precisely on a marked pad set into the stone tiles. The matte-gray frame stands out against the backdrop of greenery and pastel buildings, completing its task in one smooth motion.
C.2 Extracting Physical Phenomena from Training Dataset
Given the cost associated with using the GPT API, we utilize DeepSeek-R1-32B [14] to extract physical phenomena from large-scale training data. We provide illustrative examples and instruct DeepSeek to generate physical rules in JSON format. The prompt is shown below.
Instruction for Extracting Physical Phenomena Task: Enhance video captions by (1) identifying physical rules, then (2) rewriting the text explicitly using those rules. Goal: Generate 5 diverse, faithful rewrites of the original description, grounded in physics. Instructions: 1. Physical Rules Extraction: Analyze the input description for observable physical phenomena; Output JSON format: "Observation": "Physics Principle", and Wrap rules with <physics_rules></physics_rules>. 2. Rewriting Requirements: Expand descriptions using identified physics principles; Create 5 variants with different styles that may include (but not limited to) the following features: (a) Naturally integrate physics terms (friction, momentum, etc.) into narrative (b) Vary sentence structure complexity (concise vs elaborate) (c) Use hybrid technical/natural language (e.g. "friction (surface adhesion)" parentheses) (d) Maintain original semantic core 3. Keep faithful to original content 4. Wrap each rewrite with <rewrite1></rewrite1> through <rewrite5></rewrite5> Example Input: "A Segway bumps over small speed bumps, rider maintaining balance." Example Output: <physics_rules> {{ "The Segway wheels rotate.": ["Friction"], "The rider and Segway move forward.": ["Conservation of Momentum"], "The rider’s shadow length changes as the sun’s angle changes.": ["Reflection"], "The Segway should experience a change in speed when going over speed bumps": ["Friction"] }} </physics_rules> <rewrite1>A Segway navigates speed bumps, its rubber tires generating friction as they rotate against the pavement. The rider leans forward, conserving momentum through the obstacle course while their elongated shadow reflects the low sun angle. Each bump momentarily reduces speed as friction increases between wheels and concrete.</rewrite1> Now process: "{original_caption}" Required Output Format: <physics_rules> {{JSON}} </physics_rules> <rewrite1>…</rewrite1> <rewrite5>…</rewrite5>
Although we also prompt DeepSeek to produce various versions of enhanced prompts, we find that the original video captions generally offer higher quality. This is because DeepSeek often generates imagined scenarios that do not accurately reflect the video content. As a result, we retain the original captions for training and use only the physical rules and phenomena extracted by DeepSeek.
Figure 7: DiffPhy enables physically reasonable and semantically coherent video generation in challenging cases. (1) shows the comparison to T2V models and (2) provides multi-frame comparison to Wan 2.1-14B. DiffPhy generates more semantically coherent and physically reasonable videos compared to other methods.
C.3 Comparison of Enhanced Prompts
To demonstrate the effectiveness of our proposed CoT reasoning, we present a prompt comparison using the videos shown in Figure 7-(1) as an example. First, we show the original user input, which is a short descriptive sentence. Next, we include the enhanced prompt generated by VideoPhy, followed by the prompt produced by our Chain-of-Thought (CoT) method. For reference, we also provide the phenomena checklist used by VideoPhy. Finally, we provide discussion about the advantage of our enhanced prompts.
(a) Category distribution
(b) Video examples
Figure 8: Illustration of PhyHQ dataset. (a) shows the distribution of video categories and (b) shows examples of video from PhyHQ dataset.
Input:
A finger pokes a small stuffed animal hanging from a bedpost, causing it to fall to the floor.
Enhanced prompt by VideoPhy:
In a cozy, dimly lit bedroom, a small, plush stuffed animal dangles from a bedpost, its soft, fluffy body swaying gently in the still air. The camera captures a close-up of the toy, its vibrant blue fur and large, round eyes inviting warmth and comfort. As the scene unfolds, a human finger, adorned with a delicate ring, enters the frame, its tip gently pressing against the toy’s head. The soft pressure causes the toy to sway, its head tilting slightly, evoking a playful, interactive moment. The background remains blurred, enhancing the focus on the toy and the finger, while the lighting casts a gentle glow, creating a serene atmosphere. The finger’s movement is deliberate yet tender, as it continues to poke and nudge the toy, ultimately causing it to tumble down, landing softly on the floor. The camera remains static, capturing the subtle interplay of touch and the toy’s descent, inviting viewers into a world of playful exploration.
Enhanced prompt by DiffPhy:
A slender finger reaches out and pokes a small stuffed animal hanging from a bedpost. The plush toy wobbles and tilts, its stitches rustling softly before it loses its grip and falls to the floor. The gentle thud echoes in the quiet room, where soft light filters through curtains, casting warm patterns across the carpet.
Physical phenomena:
Elasticity: The finger makes contact with the toy, causing a slight deformation of the toy’s fur. Gravity: The toy remains attached to the shelf after being poked. Friction: The toy does not move significantly from its original position on the shelf after being poked.
Discussion:
Although VideoPhy generates rich contextual descriptions, it tends to focus more on atmospheric details rather than the physical interactions themselves. As a result, while the scenes may appear vivid, the generated videos often fail to capture key interactions—such as poking or objects falling to the floor. In contrast, the enhanced prompts produced by our CoT reasoning approach explicitly describe both the critical physical phenomena and the surrounding atmosphere. The descriptions of physical interactions ensure physical commonsense and semantic coherence, while the atmospheric details contribute to overall naturalness. We also include the physical phenomena identified by VideoPhy for reference. As shown in Figure 7, our generated video aligns well with these phenomena, demonstrating the effectiveness of our approach.
Appendix D Additional Qualitative Comparisons
For qualitative comparison, we included results against existing T2V methods in the main manuscript. Here, Figure 7 provides additional comparisons between our approach and existing methods. Specifically, Figure 7(a) highlights differences in semantic coherence when evaluated against several state-of-the-art text-to-video models, including CogVideoX, Cosmos, Hunyuan, Sora, and Wan. Figure 7(b) offers a detailed frame-by-frame comparison between Wan and our DiffPhy model across six frames. For a more comprehensive understanding and additional qualitative insights, we encourage readers to view our supplemental demo video.
Appendix E Details of PhyHQ Dataset
The PhyHQ dataset consists of 8,000 videos, each paired with a caption and corresponding physical phenomena, curated according to the procedure described in Section B. Figure 7 illustrates the category distribution and provides example videos from PhyHQ. The dataset spans a wide range of categories, including people, vehicles, food, animals, technology, urban landscapes, natural scenery, plants, oceans, architecture, and outer space. As illustration, we provide the detailed text prompts used in cases (a) and (b) below. (a) In the video, a man is seen using a sledgehammer to break down a brick wall. He is wearing a gray shirt, black shorts, and a black cap. The man is standing on a pile of rubble and debris, indicating that he has been working on this task for some time. The wall appears to be part of a larger structure, possibly a building or a house. The man is focused on his task, and his movements are deliberate and forceful as he swings the sledgehammer to break the bricks. The sound of the hammer hitting the bricks is loud and echoes in the surrounding area. The man’s body language suggests that he is exerting a significant amount of effort to complete this task. (b) The video shows a man driving a boat on a body of water. The boat is moving at a high speed, creating a large wake behind it. The man is wearing a green hat and appears to be enjoying the ride. The water is a deep blue color, and the sky is clear. The boat is white and appears to be a small motorboat. The man is the only person visible in the video.
Appendix F Limitations and Future Work
Despite recent advances in multimodal large language models (MLLMs), their ability to reason and evaluate physical visual content remains limited. While these models perform well on tasks involving textual physical reasoning, they often struggle to interpret videos, particularly in evaluating the detailed alignment between video and text, or determining the physical commonsense of complex scenarios. Current outputs from these models tend to be shallow or generic, lacking the granularity required for tasks that demand nuanced feedback. This highlights the need for more robust mechanisms to process and reason over long-sequence video data.
To address these shortcomings, we propose two key directions. First, enhancing the quality and diversity of the physical video dataset is essential. Our current dataset comprises 8,000 curated videos, but scaling this further, i.e., applying more fine-grained filtering and quality control, could significantly improve model learning. Second, future work should explore training strategies that explicitly target video-based reasoning and feedback generation, possibly through better failure cases solving mechanism. These steps could enable MLLMs to produce more detailed, context-aware assessments, ultimately closing the gap between textual and video understanding.